1Two annuities and two zero coupon bonds are available to buy or sell today. Their characteristics are: • The first annuity, AWhat is the forward price to buy the annuity A1 at time T = 2, just after the annuity payment is made?
Answer #1

For A1, to calculate the forward price of this annuity we have to first calculate the interest rate

Using this formula, 1-(1+i)-n PV = C*

PV= $8743.27 C= $2000 and n =5

i = 4.65%

Now at Time(T)= 2, we will have to calculate the FV of the bond using this formula,FV =C+(1+i) - 1

But in this case, n= 3 years since 2 years have passed

FV = 2000 * 3.1417 = 6283.32

Hence, the forward price to buy the annuity A1 would be $6283.32